# Kennedy-Tasaki transformation and non-invertible symmetry in lattice models beyond one dimension

### v1

@@ VI.1 Kramers-Wannier for SSSB Hamiltonian
For simplicity, we consider the model on a square lattice with $L_{x}=L_{y}=L$. One can think of the first term as the product of two plaquette Ising term shifted in both $x$ and $y$ directions by one. The model is symmetric under a set of unitary transformations which represents spin flips along rigid lines in $x$, $y$, and diagonal directions:
@@
$\mathbf{D}_{\text{DPIM}}\equiv\mathbf{P}_{\text{DPIM}}\tilde{\mathbf{D}}_{y}\mathbf{H}\tilde{\mathbf{D}}_{x}\mathbf{H}\tilde{\mathbf{D}}_{\text{diag}}\mathbf{P}_{\text{DPIM}},$ (77)
where $\mathbf{H}$ is the simultaneous Hadamard transformation on all the qubits, as before. In the bulk, the transformation occurs as below:
$\ \boxed{X}&&\end{array}\overset{\tilde{\mathbf{D}}_{\text{diag}}}{\longrightarrow}\begin{array}[]{ccc}&&\\ &Z&\\ \boxed{Z}&&\end{array}\overset{\mathbf{H}}{\longrightarrow}\begin{array}[]{ccc}&&\\ &X&\\ \boxed{X}&&\end{array}\overset{\tilde{\mathbf{D}}_{x}}{\longrightarrow}\begin{array}[]{ccc}&&\\ &Z&Z\\ \boxed{Z}&Z&\end{array}\overset{\mathbf{H}}{\longrightarrow}\begin{array}[]{ccc}&&\\ &X&X\\ \boxed{X}&X&\end{array}\overset{\tilde{\mathbf{D}}_{y}}{\longrightarrow}\begin{array}[]{ccc}&Z&Z\\ Z&&Z\\ \boxed{Z}&Z&\end{array}\ ,$ (96)
@@
for $\eta\in\{\eta^{x}_{j},\eta^{y}_{i},\eta^{\text{diag}}_{k}\}$; in particular at the self-duality point $\lambda=1$, we get an algebra involving a non-invertible symmetry. Moreover, with appropriate choice of phase normalization for $\mathbf{D}_{\text{DPIM}}$, we have
$\left(\mathbf{D}_{\text{DPIM}}\right)^{2}=\mathbf{T}_{(2,2)}\mathbf{P}_{\text{DPIM}},$ (98)
where $\mathbf{T}_{(2,2)}$ is the translation by $(2,2)$. Note, however, that one could set the opposite ordering for the diagonal part of the unitary. Then such a KW operator $\mathbf{D}^{\prime}_{\text{DPIM}}$ defined with it would obey $\left(\mathbf{D}^{\prime}_{\text{DPIM}}\right)^{2}=\mathbf{P}_{\text{DPIM}}$, so the translation is not an essential feature in this case.
@@ VI.2 Kennedy-Tasaki for SSSB and SSPT
Now note that the map $\mathbf{D}_{\text{DPIM}}$ preserves locality for symmetric operators (those composed of single $X$ and the product of six $Z$’s), and it also preserves a gap. The property $\eta\mathbf{D}_{\text{DPIM}}=\mathbf{D}_{\text{DPIM}}$ means that this is a gauging map [37, 38, 39, 40]. We expect that there is no finite-depth local unitary circuit connecting the symmetric ground states described by the two stabilizers on the right-hand side: the one being the cluster state (short-range entangled) and the other being SSSB with long-range order. Then one cannot have a symmetric (under $\{\eta^{x}_{j},\eta^{y}_{i},\eta^{\text{diag}}_{k}\}$) finite-depth local unitary connecting the states described by the two stabilizers on the left hand side. Since $X$ is the stabilizer for the trivial symmetric state, other one has to belong to a non-trivial SSPT.
Finally, we can define a Kennedy-Tasaki transformation:
@@ VI.3 Three dimensions and higher
$\mathbf{D}_{\text{DCIM}}\equiv\mathbf{P}_{\text{DCIM}}\tilde{\mathbf{D}}_{z}\mathbf{H}\tilde{\mathbf{D}}_{y}\mathbf{H}\tilde{\mathbf{D}}_{x}\mathbf{H}\tilde{\mathbf{D}}_{\text{diag}}\mathbf{P}_{\text{DCIM}},$ (121)
with an appropriate definition of $\tilde{\mathbf{D}}_{\text{diag}}$ and $\mathbf{P}_{\text{DCIM}}$, where the latter imposes all the symmetry generators $\{\eta^{x}_{j},\eta^{y}_{i},\eta^{z}_{m},\eta^{\text{diag}}_{k}\}$ to be evaluated to unity for the ground states. There is also a cluster state whose stabilizer is given by one $X$ (sitting at the overlapping site of two diagonally neighboring cubes) and fourteen $Z$’s (sitting at the remaining corners), which is produced by a combination of $CZ$ gates; see Fig. 6(b). The cluster-state entangler and the Kramers-Wannier duality operator $\mathbf{D}_{\text{DCIM}}$ form a web of dualities.
@@ VII Conclusion
In this paper, we have presented a higher dimensional generalization of the non-invertible Kramers-Wannier duality symmetry on a lattice. The generalized hypercubic Ising models with a transverse field exhibit non-invertible symmetry at the self-dual point. In addition to that, we have also presented a generalization of the Kennedy-Tasaki transformation in higher dimensions. In our examples involving $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT phase, under the KT transformation, the higher dimensional cluster-state model with an external field decomposes into two copies of hypercubic transverse-field Ising models. The KT transformation derived in the main text is obtained by sandwiching the cluster-state entangler between the KW duality operator (on two sublattices) and its Hermitian conjugate. We have also derived an alternative KT transformation in the appendix B. Both of these variants of the KT transformation achieve the effect of taking a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ model to two copies of SSSB models. Our result generalizes the picture of the 1D KT transformation proposed by [25] to higher dimensions. In addition to the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry, we also discussed KW duality symmetry of DHCIM and used that to give a KT transformation that maps between the cluster model, which is $\mathbb{Z}_{2}$ SSPT, and one copy of DHCIM that is in the SSSB phase.
@@
While the KW operators in our construction require linear-depth circuits, they can also be implemented using a cluster-state entangler acting on the original and ancillary degrees of freedom, measurement on the original degrees of freedom, and then a feedforward correction, the whole combination of which is finite-depth. It is, therefore, feasible to implement the KW and KT transformations on quantum devices, following the approaches in, e.g., Refs. [41, 42, 43, 44]. On the other hand, in the case of 1D [16], the linear-depth construction of KW duality operator (on the same lattice) is a non-invertible symmetry arising from the anomalous translation symmetry in the Majorana fermion representation after gauging the global fermion parity. Hence, this construction clarifies the relationship between anomalies and non-invertible symmetry. In appendix. A, we also discuss Majorana hypercube models and identify the exchange in the Majorana terms that gives rise to the KW duality in the corresponding Ising hypercube model. The exchange action is not related to translation and, moreover, does not commute with the subsystem fermion parity. However, the physical meaning of this exchange symmetry is not yet clear.
@@
The non-invertible symmetries would also have a generalization to a Kramers-Wannier transformation that gauges higher-form symmetries. In Ref. [45], for example, gauging $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1-form symmetry was considered using a mathematical map, which transforms the 3D cluster state (also called the Raussendorf-Bravyi-Harrington state [46]) to itself and the product state to two copies of the 3D toric code; essentially, this is a higher-form symmetry generalization of the Kramers-Wannier transformation. We have confirmed that the parent Hamiltonians of the above mentioned states map in the same corresponding way, under the measurement-assisted construction. It would be interesting to explore the KW transformation more broadly beyond this example. One could also define a higher-form generalization of the Kennedy-Tasaki transformation by composing the map with the 3D cluster state entangler, which in total brings an SPT state to some copies of SSB states with respect to the higher-form symmetry. It would be interesting to construct a circuit with a projector that realizes this transformation and study its algebra.
@@
It is also natural to study the Kennedy-Tasaki transformations that map between SSPT phase and SSSB phase in higher-dimensional models for more examples and other symmetry groups. Classification of SSPT phases beyond two dimensions is not known yet. Thus, it would be worthwhile to extend the classification of SSPT phases to higher dimensions and find a KT transformation that maps between all the SSPT to SSSB phases for symmetry groups beyond $\mathbb{Z}_{2}$ and $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
@@
The existence of non-invertible symmetry put constraints similar to the Lieb-Shultz-Mattis theorem on the low energy theory of one-dimensional lattice models: it is recently found in the case of 1D by Seiberg, Seifnashri and Shao [47] that the system is either in a gapless phase or gapped phase with a three (or a multiple of three) degenerate ground states at the non-invertible symmetric point. It would be interesting to study similar constraints on the higher-dimensional lattice models with the non-invertible symmetries we have discussed in this work.

### v2

@@ VI.1 Kramers-Wannier for SSSB Hamiltonian
For simplicity, we consider the model on a square lattice with $L_{x}=L_{y}=L$. One can think of the first term as the product of two plaquette Ising terms shifted in both $x$ and $y$ directions by one. The model is symmetric under a set of unitary transformations that represents spin flips along rigid lines in $x$, $y$, and diagonal directions:
@@
$\mathbf{D}_{\text{DPIM}}\equiv\mathbf{P}_{\text{DPIM}}\tilde{\mathbf{D}}_{y}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}}\tilde{\mathbf{D}}_{x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}}\tilde{\mathbf{D}}_{\text{diag}}\mathbf{P}_{\text{DPIM}},$ (79)
where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}}$ is the simultaneous Hadamard transformation on all the qubits, as before. In the bulk, the transformation occurs as below:
$\ \boxed{X}&&\end{array}\overset{\tilde{\mathbf{D}}_{\text{diag}}}{\longrightarrow}\begin{array}[]{ccc}&&\\ &Z&\\ \boxed{Z}&&\end{array}\overset{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}}}{\longrightarrow}\begin{array}[]{ccc}&&\\ &X&\\ \boxed{X}&&\end{array}\overset{\tilde{\mathbf{D}}_{x}}{\longrightarrow}\begin{array}[]{ccc}&&\\ &Z&Z\\ \boxed{Z}&Z&\end{array}\overset{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(2)}}}{\longrightarrow}\begin{array}[]{ccc}&&\\ &X&X\\ \boxed{X}&X&\end{array}\overset{\tilde{\mathbf{D}}_{y}}{\longrightarrow}\begin{array}[]{ccc}&Z&Z\\ Z&&Z\\ \boxed{Z}&Z&\end{array}\ ,$ (98)
@@
for $\eta\in\{\eta^{x}_{j},\eta^{y}_{i},\eta^{\text{diag}}_{k}\}$; in particular at the self-duality point $\lambda=1$, we get an algebra involving a non-invertible symmetry. Moreover, $\mathbf{D}_{\text{DPIM}}$ squares to a diagonal translation by two units and a projection up to an overall normalization
$\left(\mathbf{D}_{\text{DPIM}}\right)^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\propto}\,\mathbf{T}_{(2,2)}\mathbf{P}_{\text{DPIM}},$ (100)
where $\mathbf{T}_{(2,2)}$ is the translation by $(2,2)$. Note, however, that one could set the opposite ordering for the diagonal part of the unitary. Then such a KW operator $\mathbf{D}^{\prime}_{\text{DPIM}}$ defined with it would obey $\left(\mathbf{D}^{\prime}_{\text{DPIM}}\right)^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\propto}\,\mathbf{P}_{\text{DPIM}}$, so the translation is not an essential feature in this case. We also comment that we can implement the KW transformation in a short-depth operation (including measurement) in appendix F.
@@ VI.2 Kennedy-Tasaki for SSSB and SSPT
Now note that the map $\mathbf{D}_{\text{DPIM}}$ preserves locality for symmetric operators (those composed of single $X$ and the product of six $Z$’s), and it also preserves a gap. The property $\eta\mathbf{D}_{\text{DPIM}}=\mathbf{D}_{\text{DPIM}}$ means that this is a gauging map [43, 44, 45, 46]. We expect that there is no finite-depth local unitary circuit connecting the symmetric ground states described by the two stabilizers on the right-hand side: the one being the cluster state (short-range entangled) and the other being SSSB with long-range order. Then one cannot have a symmetric (under $\{\eta^{x}_{j},\eta^{y}_{i},\eta^{\text{diag}}_{k}\}$) finite-depth local unitary connecting the states described by the two stabilizers on the left-hand side. Since $X$ is the stabilizer for the trivial symmetric state, the other one has to belong to a non-trivial SSPT. Finally, we can define a Kennedy-Tasaki transformation:
@@ VI.3 Three dimensions and higher
$\begin{split}\mathbf{D}_{\text{DCIM}}&\equiv\mathbf{P}_{\text{DCIM}}\tilde{\mathbf{D}}_{z}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(3)}}\tilde{\mathbf{D}}_{y}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(3)}}\times\\ &\qquad\qquad\tilde{\mathbf{D}}_{x}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(3)}}\tilde{\mathbf{D}}_{\text{diag}}\mathbf{P}_{\text{DCIM}},\end{split}$ (123)
with an appropriate definition of $\tilde{\mathbf{D}}_{\text{diag}}$ , ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{H}^{\otimes(3)}}$ and $\mathbf{P}_{\text{DCIM}}$, where the latter imposes all the symmetry generators $\{\eta^{x}_{j},\eta^{y}_{i},\eta^{z}_{m},\eta^{\text{diag}}_{k}\}$ to be evaluated to unity for the ground states.
On the other side of the duality web, there is also a cluster state whose stabilizer is given by one $X$ (sitting at the overlapping site of two diagonally neighboring cubes) and fourteen $Z$’s (sitting at the remaining corners), which is produced by a combination of $CZ$ gates; see Fig. 6(b). This 3D cluster state is a nontrivial 3D $\mathbb{Z}_{2}$ SSPT; see Ref. [47] for other generators of 3D $\mathbb{Z}_{2}$ SSPT phases. The cluster-state entangler and the Kramers-Wannier duality operator $\mathbf{D}_{\text{DCIM}}$ form a web of dualities.
@@ VII Conclusion
In this paper, we have presented a higher-dimensional generalization of the non-invertible Kramers-Wannier duality symmetry on a lattice. The generalized hypercubic Ising models with a transverse field exhibit non-invertible symmetry at the self-dual point. In addition to that, we have also presented a generalization of the Kennedy-Tasaki transformation in higher dimensions. In our examples involving $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ SSPT phases, under the KT transformation, the higher-dimensional cluster-state model with an external field decomposes into two copies of hypercubic transverse-field Ising models. The KT transformation derived in the main text is obtained by sandwiching the cluster-state entangler between the KW duality operator (on two sublattices) and its Hermitian conjugate. We have also derived an alternative KT transformation in the appendix B. Both of these variants of the KT transformation achieve the effect of taking a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ model to two copies of SSSB models. Our result generalizes the picture of the 1D KT transformation proposed by [27] to higher dimensions. In addition to the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry, we also discussed KW duality symmetry of DHCIM and used that to give a KT transformation that maps between the cluster model, which is $\mathbb{Z}_{2}$ SSPT, and one copy of DHCIM, which is in the SSSB phase.
@@
While the KW operators in our construction require linear-depth circuits, they can also be implemented using a cluster-state entangler acting on the original and ancillary degrees of freedom, measurement on the original degrees of freedom, and then a feedforward correction, the whole combination of which is finite-depth (see appendix F). It is, therefore, feasible to implement the KW and KT transformations on quantum devices, following the approaches in, e.g., Refs. [48, 49, 50, 51]. On the other hand, in the case of 1D [18], the linear-depth construction of the KW duality operator (on the same lattice) is a non-invertible symmetry arising from the anomalous translation symmetry in the Majorana fermion representation after gauging the global fermion parity. Hence, this construction clarifies the relationship between anomalies and non-invertible symmetry. In appendix. A, we also discuss Majorana hypercube models and identify the exchange in the Majorana terms that gives rise to the KW duality in the corresponding Ising hypercube model. The exchange action is not related to translation and, moreover, does not commute with the subsystem fermion parity. However, the physical meaning of this exchange symmetry is not yet clear.
@@
The non-invertible symmetries would also have a generalization to a Kramers-Wannier transformation that gauges higher-form symmetries. In Ref. [52], for example, gauging $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1-form symmetry was considered using a mathematical map, which transforms the 3D cluster state (also called the Raussendorf-Bravyi-Harrington state [53]) to itself and the product state to two copies of the 3D toric code; essentially, this is a higher-form symmetry generalization of the Kramers-Wannier transformation. We have confirmed that the parent Hamiltonians of the above-mentioned states map in the same corresponding way under the measurement-assisted construction. It would be interesting to explore the KW transformation more broadly beyond this example. One could also define a higher-form generalization of the Kennedy-Tasaki transformation by composing the map with the 3D cluster state entangler, which in total brings an SPT state to some copies of SSB states with respect to the higher-form symmetry. It would be interesting to construct a circuit with a projector that realizes this transformation and study its algebra.
@@
It is also natural to study the Kennedy-Tasaki transformations that map between the SSPT phase and the SSB phase in higher-dimensional models for more examples and other symmetry groups. Classification of SSPT phases beyond two dimensions is not extensively known yet. Thus, it would be worthwhile to extend the classification of SSPT phases to higher dimensions and find a KT transformation that maps between all the SSPT to SSSB phases for symmetry groups beyond $\mathbb{Z}_{2}$ and $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$.
Recently, KT has been used to discuss SPT phases protected by non-invertible symmetry [54] in 1+1 dimensions. The distinct non-invertible SPT phases are characterized by various symmetry-breaking patterns of the dual symmetry that emerges after applying the KT transformation. It would be interesting to analyze such non-invertible SSPT phases in 2+1 dimensions. We speculate that the cluster states with $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ or $\mathbb{Z}_{2}$ subsystem symmetry would split into distinct SSPTs after imposing the non-invertible KW duality symmetry. However, we leave a comprehensive analysis for future exploration.
@@
The existence of non-invertible symmetry put constraints similar to the Lieb-Shultz-Mattis theorem on the low energy theory of one-dimensional lattice models: it is recently found in the case of 1D by Seiberg, Seifnashri, and Shao [55] that the system is either in a gapless phase or gapped phase with a three (or a multiple of three) degenerate ground states at the non-invertible symmetric point. Studying such LSM constraints with non-invertible symmetries beyond one dimension would be generally interesting. In two dimensions, the immediate question is whether lattice models with $\mathbb{Z}_{2}$ subsystem symmetry and the non-invertible symmetry $\mathbf{D}^{(2)}$ obey any such constraints. Let us consider some possible $\mathbb{Z}_{2}$ subsystem symmetric phases on the two-dimensional torus.
  • According to the classification of strong SSPTs in two dimensions, there are
    • the trivial SSPT ($\rm H_{1}$ in the first row),
    • one non-trivial $\mathbb{Z}_{2}$ SSPT($\rm H_{1}$ in the second row).
  • There is an SSSB phase ($\rm H_{2}$ in the first row).
  • There is a topological order, i.e., the Wen-plaquette model [56] ($\rm H_{2}$ in the second row).
The trivial SSPT is mapped to the SSSB phase under the Kramers-Wannier operator $\mathbf{D}^{(2)}$; see the first row of table 2. It is easy to see that the non-trivial $\mathbb{Z}_{2}$ SSPT that was discussed in section VI.2 is mapped to the Wen-plaquette model under $\mathbf{D}^{(2)}$; see the second row of table 2. The trivial and non-trivial SSPTs have a single ground state. The SSSB phase has $2^{L_{x}+L_{y}-1}$ ground states and this can be seen by counting the number of independent stabilizer generators. The topological order (the Wen-plaquette model) has a four or two-fold ground state degeneracy depending on whether it is an even by even lattice or not [56]. At the critical point between the trivial phase and SSSB phase (first-order transition), we expect $12^{L_{x}+L_{y}-1}+$ 1 ground states. The critical point between non-trivial SSPT and topological order is either gapless or has five or three ground states depending on the lattice size. We summarize our discussions so far about ground state degeneracy of various models and their KW dual that are $\mathbb{Z}_{2}$ subsystem symmetric in table 2 with two more examples.
$\rm H_{1}$ term $\rm H_{2}$ term GSD of $\rm H_{1}$ GSD of $\rm H_{2}$
$\begin{array}[]{cc}&X\\ &\end{array}$ $\begin{array}[]{cc}Z&Z\\ Z&Z\end{array}$ 1 $2^{L_{x}+L_{y}-1}$
$\begin{array}[]{ccc}&Z&Z\\ Z&X&Z\\ Z&Z&\end{array}$ $\begin{array}[]{ccc}&Z&Y\\ &Y&Z\\ &&\end{array}$ 1
4 for $L_{x}$, $L_{y}$ even,
2 otherwise
$\begin{array}[]{ccc}&X&X\\ &&\end{array}$ $\begin{array}[]{ccc}Z&I&Z\\ Z&I&Z\end{array}$ $2^{L_{y}}$
$2^{L_{x}+2L_{y}-2}$ for $L_{x}$ even,
$2^{L_{x}+L_{y}-1}$ for $L_{x}$ odd
$\begin{array}[]{ccc}&&X\\ &X&\\ &&\end{array}$ $\begin{array}[]{ccc}&Z&Z\\ Z&&Z\\ Z&Z&\end{array}$
$2^{L}$
for $L_{x}=L_{y}=L$
$2^{3L-2}$
for $L_{x}=L_{y}=L$
Table 2: Ground state degeneracy of various Hamiltonians with $\mathbb{Z}_{2}$ subsystem symmetry that we denote by $\rm H_{1}$ and $\rm H_{2}$. We omit the summation over lattice sites and overall negative signs in front of the Hamiltonian terms for simplicity. $\rm H_{1}$ and $\rm H_{2}$ are related by a Kramers-Wannier duality, i.e., $\rm\mathbf{D}^{(2)}\rm H_{1}=H_{2}\mathbf{D}^{(2)}$.
For more general Hamiltonians that are invariant under both $\mathbb{Z}_{2}$ subsystem symmetry and non-invertible symmetry $\mathbf{D}^{(2)}$, counting the number of ground states requires a more careful analysis. This is because, first of all, the constraint about the number of ground states in the spirit of the LSM theorem is a question about the thermodynamic behavior of the lattice models that we start with. For lattice models with subsystem symmetries, the thermodynamic limit is very sensitive to the lattice system size. This is manifest in the extensive ground state degeneracy of these lattice models with large but finite system sizes. It would be interesting if one could define a notion of ground-state degeneracy in this setup as well as in higher dimensions and find an LSM-type constraint on the number of ground states. However, we leave this question for future exploration.